We have decided that the IFO alignment is bad enough that we're not ready to pump down. PUMP ABORTED.
The IFO alignment is somewhat OK, in that the green and IR beams are flashing in the arms, and the return beams are overlapping at the BS. However the beams appear to be not centered on any of the optics at the moment. They are all displaced in yaw by ~0.5 to 1 cm or so in various directions.
From this we have decided that we need to step back and reattack the IFO alignment from square one. Here is our current suggested procedure:
- check ETM positions relative to what we think they should be on the drawings. This is to verify that the ETMs were not placed in the wrong places laterally.
- translate Y green axis north, centering green on ETMY and ITMY (by looking at cards). North is the opposite direction from how the beams are displaced from the TM centers.
- steer input pointing to overlap IR on green beam at BS, ITMY, and ETMY. IR should visibly overlap green at both BS and ITMY, and we should be able to see IR on target in front of ETMY with ETMY face camera, and in ETMY trans camera.
- center IR on ETMX by steering BS with DC bias.
- align Y arm cavity for green resonance by adjusting ITMY.
- adjust ITMX to achieve michelson fringes at AS
- adjust PRM lateral translation to center beam on PRM, if needed
- adjust SRM lateral translation to center beam on SRM, if needed
- align PRC to see fringes
- align SRC to see fringes
- extract AS (no clipping)
Once this is done, we will need to check the following:
- IPANG/IPPOS extraction
- pick-off extraction
- green periscopes and green beam extraction at PSL
We've decided to stop for the night, get a good nights rest, and attack all of this tomorrow morning.